OpenStudy (anonymous):

Simplify using the order of operations: (5+3)3 – 10/2=

OpenStudy (anonymous):

is it (5+3)^3 or (5+3)*3?

OpenStudy (anonymous):

(5+3)*3?

OpenStudy (anonymous):

=8*3-5 -24-5 =19

OpenStudy (amistre64):

(5+3)3 – 10/2 3(5)+3(3) – 10/2 15+9 – 10/2 (15(2)+9(2) – 10)/2 (30+18–10)/2 (20+18)/2 38/2 = 15+4 = 19